OpenStudy (anonymous):

How do I find net momentum? In the problem I am given that there is a 2300 kg truck traveling at 19 m/s in the -x direction, and a 1600 kg car traveling in the positive y direction. How do I find net momentum, and the direction of the net momentum?

OpenStudy (anonymous):

@mathslover, could you help me with this?

mathslover (mathslover):

What is the velocity of the 1600 kg car

mathslover (mathslover):

Is it 19 m/s

OpenStudy (anonymous):

I'm sorry, i meant to add it. It is 25 m/s

mathslover (mathslover):

Ok

mathslover (mathslover):

Now, I am going to use the law of vector addition

mathslover (mathslover):

See, Let the car having mass of 2300 kg = A and that having 1600 kg = B

mathslover (mathslover):

Now, total linear momentum, P net vector = P (A) vector + P (B) vector

mathslover (mathslover):

Now what is momentum ?

mathslover (mathslover):

I mean what is momentum in terms of mass and velocity

mathslover (mathslover):

@dave0616

OpenStudy (anonymous):

momentum = p=m*v. The thing is though that I typed that into my online homework and it is not accepting it for some reason. I just want to make sure i am doing it right. I would assume that it is not looking for a signed value though, because it asks for the magnitude

mathslover (mathslover):

I accept that

mathslover (mathslover):

Now, What is the momentum of A ? It is, P(A) vector = -2300 * 19 i cap

mathslover (mathslover):

And P (B) vector = 1600 * 25 j cap

mathslover (mathslover):

Got it

OpenStudy (anonymous):

so would you just add the value of those two together? P(total)= (-2300*19)+(1600*25)=-3700?

mathslover (mathslover):

No dude

mathslover (mathslover):

You get total momentum as P vector = P (A) vector + P (B) vector P (vector) = -2300 * 10 i cap + 1600 * 25 j cap Now, magnitude of momentum will be \[P = \sqrt{(-2300 * 10)^2 + (1600*25)^2}\]

mathslover (mathslover):

This will be the magnitude of total linear momentum

mathslover (mathslover):

Got it ?

OpenStudy (anonymous):

that makes complete sense, thank you very much mathslover

mathslover (mathslover):

Your welcome

OpenStudy (anonymous):

how would i go about finding the direction of the total momentum?

mathslover (mathslover):

Allright Lets make a nice diagram for that

mathslover (mathslover):